OpenStudy (anonymous):

What is the simplified form of the quantity ysquared minus 6y plus 8 over the quantity ysquared minus 7y plus 12 ? HELP

OpenStudy (anonymous):

\[(y^2 - 6y + 8)/(y^2 -7y + 12)\] ?

OpenStudy (anonymous):

I mean is that what you're asking?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(y-2)(y-4) / (y-3)(y-4)

OpenStudy (anonymous):

Numerator: -two numbers that multiply to 8 and add to -6: -4, -2 Therefore, (y^2 - 6y + 8) = (y - 4)(y - 2) Denominator: -two numbers that multiply to 12 and add to -7: -4, -3 Therefore, (y^2 - 7y + 12) = (y - 4)(y - 3) So the new fraction is: \[[(y - 4)(y - 2)]/ [(y - 4)(y - 3)]\] Then because it is all multiplication and division you can cancel the (y - 4)s: \[(y - 2)/(y - 3)\]

OpenStudy (anonymous):

the (y-4) cancel and you're left with (y-2)/(y-3)

OpenStudy (anonymous):

THANKS GUYS :)

OpenStudy (anonymous):

you're welcome. makes sense?

OpenStudy (anonymous):

yup